# Evaluation of ChatGPT Usability as A Code Generation Tool

### v2

@@ IV-C Difficulty Levels
Moreover, when measuring usability on each individual quality criterion, the experiment data also clearly demonstrates the monotonicity of performance with difficulty level; see Figure 13.
@@
Figure 13: Usability on Various Quality Criteria for Tasks of Different Difficulty Levels.
IV-D Learnability
The use of LLMs like ChatGPT to generate program code with natural language input is a process of interactions between the human user and the ML system. It has been widely recognized that the natural language input plays a significant part to the success of generating high quality outputs. This leads to a research question whether it is easy for the human user to learn how to make good input to the ML system. To answer this question, we conducted the testing and evaluation in a sequential process from test set $T_{0}$ to $T_{4}$ and analyzed the differences between the results.
Because the test cases in various test sets are drawn at random from a large dataset, the questions in different test sets have different levels of difficulty. In order to take this factor into consideration, we calculate the difficulty levels of the test sets using the following formula.
$DL(T)=(1\times E_{T}+2\times M_{T}+3\times H_{T})/(E_{T}+M_{T}+H_{T})$
where $DL(T)$ denotes the difficulty level of a test set $T$, $E_{T}$, $M_{T}$ and $H_{T}$ are the numbers of easy, medium, and hard questions in test dataset $T$, respectively. Figure 14 shows the difficulty levels of the test sets.
Figure 14: Difficulty Levels of the Test Sets.
The quality scores of various test sets are then adjusted according to the difficulty levels; see Figure 15 and 16 for the results. From the data, we cannot observe any improvement in the quality when testing progresses from $T_{0}$ to $T_{4}$;
Figure 15: Quality Scores of Test Sets On Various Quality Attributes After Adjusted by Difficulty Levels.
Figure 16: Total Quality Scores of Test Sets After Adjusted By Difficulty Levels.
It is also observed that even the average numbers of attempts and the average completion times do not improve with the increases of experiences in using ChatGPT; see Figure 17 and 18.
Figure 17: Average Number of Attempts on Various Test Sets at Different Difficulty Levels.
Figure 18: Average Completion Time (Second) on Various Test Sets at Different Difficulty Levels.
Figure 19: Average Quality Scores on Tasks of Difficulty Levels at Different Difficulty Levels.
Figure 19 contains three charts that show the performance changes over 5 test sets on various quality attributes for the easy, medium, and hard code generation tasks, respectively. None of these three charts shows an obvious quality improvement trend with the increase of experiences in using ChatGPT.
Figure 20 shows the changes in the average total quality scores on tasks of three different difficulty levels. It also shows a lack of improvement with the increase of experiences.
Figure 20: Total Quality Scores of Test Sets at Different Difficulty Levels.
Therefore, we conclude that it is not easy to learn how to use ChatGPT to generate program code through experiences of using the ML model.
V Conclusion
V-A Summary and Comparison with Existing Works
Existing evaluations of LLM models are mostly conducted through benchmarking on standard test datasets. While this enables objective comparisons of LLMs, the results hardly reflect their usability from the user’s perspective. This is particularly important when deciding on whether to use a LLM in a specific use scenario. The difficulty is that there is a wide range potential uses of LLMs, while testing and evaluating a LLM is expensive and difficult. This paper proposed a new method that aims at assessing and evaluating LLM’s performance to reflect its usability in various use scenarios efficiently and effectively. The method consists of the following elements.
First, metadata that describe the usages of the test cases are assigned to each test case in the test dataset. This enables the flexibility of tailoring the dataset to generate subsets that represent various use scenarios, and to adjust the evaluation of the performances on different scenarios with existing test results without re-run test cases. It is different from existing datasets for evaluating code generation capability, which only contain the task (also called the problem to be solved in the literature), reference solution(s) to the problem, and a set of test data for checking the correctness of the machine generated solution.
Second, the testing on each test case takes a process that mimics the real use of the LLM. In particular, we proposed a multiple attempt process for testing LLMs. In this process, the tester tries to complete the testing task through a series of attempts until a satisfactory solution is obtained or gives up after a fixed number of maximal attempts. In each attempt, the tester formulates an input based on the previous results and the task to be completed, invokes the ML model, and then inspects the output to decide whether the solution is satisfactory. This contrasts with the single attempt process currently used by the existing work on evaluating code generation, in which a number of solutions are generated for the same input. Our process is closer to the real uses of LLMs.
Third, in our method, the quality of an output is measured by a set of metrics on usefulness from the users’ perspective. This contrasts with existing evaluations of code generation, where the functional equivalence of a machine produced solution to the reference solution(s) are used as the sole judgement on the quality of the code generated. In the real uses of LLMs for code generation, a piece of code is considered as useful even if it is not functionally correct. In many cases, a skeleton of code is acceptable even if it cannot be compiled. Our metrics focus on the usability of a solution via considering its readability, logic clarity, well-structuredness, etc. Thus, the method is closer to the real uses of LLMs.
Finally, we define two metrics on user experiences in the uses of the LLM: (1) #$\#attempt_{k}$ , i.e. the average number of attempts, and (2) the average amount of time taken to complete the test tasks. They differ from the $pass@k$ metric, which is currently used in evaluation of code generation capability. It is the probability of generating at least one correct solution when the LLM is asked to generate $k>0$ solutions. In contrast, our metrics reflect the expected number of attempts that the user will make to find a satisfactory solution, and the expected time that a user will spend on to complete a code generation task. Thus, it better represents the user experiences in the use of the LLM.
Therefore, our proposed approach to the evaluation of LLMs capability of code generation can be characterized as user centric for its focus on user experiences in the uses of the LLM. This is reflected in all aspects of the testing and evaluation, including the information contained in the test dataset, the testing process, the measurement of solution quality and the performance metrics.
Our experiment results are consistent with the good user experiences and general opinions on ChatGPT. We have also identified the weakness as well as the strength of the LLM model. In particular, it is the first time that the learnability of using ChatGPT to complete code generation tasks is studied. Our experiment data show that skills of using ChatGPT for code generation would not come naturally through experiences.
V-B Potential Threats to Validity and Future Work
@@
The scale of our experiments is small in terms of the number of subjects involved in the testing process due to our resource limitation. This forms a potential threat to the validity and generalizability of the results on the performance of ChatGPT’s code generation capability. However, the feasibility of the proposed testing method is not affected. The potential threat can be eliminated by repeated experiments with more subjects. This suggests that a topic for future work is to repeat the experiments with more subjects.
@@
For some quality metrics like readability and logic clarity, the values of metrics assigned to each test case are subjective, which depends on the subject’s knowledge, experience, and ability. This causes another potential threat to the validity of the results. Efforts have already been made to confine the threat by applying the metrics consistently according to the clear definitions of the metric values. Thus, the risk is minimized. Moreover, our experiment data show that the results obtained from such subjective judgements are consistent with objective metrics of the number of attempts and time to completion. Thus, the results should be trustable. For future research, we are working on using more objective metrics about the quality of program code developed in software engineering research, like cyclomatic complexity, coupling and cohesion, code smell, etc., in complementary to the subjective metrics proposed and used in this paper. It is worth further research to find out if machine generated codes are of high quality on these attributes designed for traditional human written code, and whether such metrics can provide meaningful assessment of the usability of machine generated code.
@@
A factor that has an impact on the experiment results is the criterion to determine whether a solution is satisfactory in the multi-attempt testing process. In principle, being user centric, it must be judged by the human users who engage in the testing. In our experiment, the guideline given to the tester is that the generated code is satisfactory if it is correct when compared with the reference solution or is close enough to the reference solution. In other words, a functionally incorrect solution may well be regarded as satisfactory if a minor editing of the machine generated solution will work. However, in practice, such reference solutions are usually unknown. Therefore, there is a gap between the judgements of satisfactoriness by the testers and the real users. This is a potential threat to the validity of the experiment results. It is desirable to give a more rigorous criterion to judge the satisfactoriness of solutions generated by machine. However, defining such a rigorous criterion is a non-trivial issue and worth further research. Study of metrics on the usefulness of machine generated solutions could be helpful.
@@
Our dataset is constructed by extracting exercise questions from textbooks. An advantage is that the questions are well presented in terms of the clarity of the language and coverage of the programming language facilities and usages as well as the correctness of the reference solution. Thus, little manual editing and filtering of data is required while other datasets constructed by extracting data from open-source code repositories have required heavy workload on manual editing and cleaning of the data. It also minimizes the possibility of testing LLM on incorrect questions. Moreover, we have managed to construct a relatively large dataset with a small amount of human resource. A potential threat to the validity of the experiment is that questions from textbooks may not represent the coding tasks in the real uses of LLMs. It is worth further research on the validity of the dataset via studying the differences between real input from users and the data in our dataset.
@@
We have included three types of metadata to our benchmark data: the type of code to be generated, the difficulty level of the task to be completed, and the source of the data. The first two represent two kinds of scenarios, which can be combined to form various specific usage scenarios as demonstrated in the experiments. Since the metadata are manually assigned to the data in the dataset, a potential threat to the validity of the experiment results can be caused by errors in the metadata. In the experiment, we have taken mitigation means to minimize the error rate by giving clear definitions of the values of metadata. For example, a coding task is assigned with “hard” as its difficulty level, if and only if “when the program code to be generated is more complicated than ten lines and requires more complex data structures and algorithms.” In the literature, some datasets do classify the coding tasks into different difficulty levels. For example, APPS dataset has three subsets of data on three difficult levels: Introductory, Interview, and Competition. However, there is no definition of the criterion on how the tasks were classified [24]. Our approach enables metadata to be assigned systematically and consistently, thus the chances of human errors are reduced. As for future work, it is worth studying other types of metadata and other values of metadata, for example, difficulty levels could be more than three. Another topic for future work is how to check the correctness of metadata.
@@
Addition to the above directions of further research that address the potential threats to the validity of the work and the generalizability of the results, there are also several other interesting topics for future work. For example, the user centric testing method proposed in this paper is expensive to conduct because of human users’ involvement in the testing process. It is worth further research to reduce the cost and to improve efficiency through testing tools and automation.
@@
It is worth noting that evaluating LLM is inevitably subjective due to the nature of NLP. Instead of trying to avoid subjective assessment, we minimize the potential bias of the assessment by decomposing the overall subject opinions into a set of quality attributes that represent different aspects of subjective opinions and use multiple levels of scores to normalize the assessments. This approach to the evaluation of software quality has been well developed and widely applied in software engineering and HCI. How to further improve the methodological rigor of user centric evaluations of LLMs will be of great importance.
@@
Finally, LLMs’ capability of generating program code is not only applicable to software development as our evaluation results show, but it is also potentially useful for other purposes such as education. For example, a student could learn coding with a LLM by raising questions in natural language and getting answers from the LLM with program codes and detailed explanations. Another possibility is to help students to understand example program code by inputting a piece of code to a LLM and getting an explanation of the code. Moreover, a student’s answer to an exercise or examination question could be input to a LLM for checking its correctness and producing a mark and feedback to the student. If a student’s answer on an exercise question is incorrect, the LLM could give the reasons why it is incorrect and delineate on how to correct it. It is worth further research to assess and evaluate the usability of LLM in such use scenarios.
@@ Acknowledgment
Appendix A. Evaluation Criteria of Quality Attributes
This appendix gives the details of the evaluation criteria for each of the quality attributes.
(1) Accuracy: The correctness and accuracy of the LLM model’s response in comparison with the standard answer to the programming question. It is assessed through executing the code produced by ChatGPT and the standard answer to the coding question. Their outputs are compared to see if the generated code gives correct output. The scores are:
1: The response is completely wrong.
2: The response contains many errors.
3: The response is reasonably accurate, shows somewhat understanding of the question but it contains some errors.
4: The response is mostly accurate but it contains a small number of minor errors.
5: The response is clear, accurate, and solves the problem comprehensively without any error.
(2) Completeness: the completeness of the solution that covers the requirements of the programming question. It is assessed by comparing the functionality of the code generated against the coding problem. The scores are:
1: The response is incomplete and addresses the question inadequately.
2: The response has significant omissions.
3: The response covers the requirements well with some minor omissions.
4: The response covers almost all requirements adequately with a small number of omissions.
5: The response is thorough, comprehensive, and covers all requirements areas.
(3) Well structuredness: the structure of the code generated and formatted. It is assessed by inspection of the code on how well the code is structured and organized. The scores are:
1: Response is poorly organized and lacks proper structure.
2: Response is somewhat organized but there are issues with the structure.
3: Response is organized well, but there is a large room for improvement for clearer structure.
4: Response is well organized and provides a clear and logical structure, however there is some room for improvement.
5: Response is highly organized and has a clear and logical structure.
(4) Conciseness: The conciseness of the code produced by ChatGPT without compromising functionality. This is assessed by review of the code in comparison with the standard answer. Typically, the generated code and standard answer are compared on the number of lines of code. The scores are:
1: The response includes significant redundancies and has a larger number of lines of code in comparison with the code of the standard answer.
2: The response has noticeable redundancies and has a notably larger number of lines compared to standard answers.
3: The response is somewhat concise but can be further reduced.
4: The response is concise with minimal redundancies.
5: The response is highly concise, conveying the in-formation efficiently, and matches the number of lines of code compared to the standard answer.
(5) Logic clarity: the coherence, soundness, and clarity of the generated code. This is examined and assessed via review of the code to identify inconsistencies or logic errors.
1: The response lacks logic and has no clear flow of thought as demonstrated in the code.
2: The response demonstrates some flow of thought, but it contains many lines of code misplaced.
3: The response demonstrates a flow of thought, but some lines of code are misplaced.
4: The response demonstrates a clear logic flow of thought but could be improved by minor rearrangement of the code.
5: The response demonstrates a clear flow of thought, and the code is logically well ordered.
(6) Parameter coverage: how well the parameters provided in the coding question are used or covered in the solution produced by the LLM model. Here, “parameter” refers to the coding elements given in the question like dataset name, variable name, data values, function names, library code names, etc. This is assessed by inspection of the code to measure if all coding elements in the question were used and appropriately incorporated in the solution.
1: The response does not use any of the parameters from the question.
2: The response uses a small proportion of parameters, however some key parameters are missed or incorrectly used.
3: The response uses some parameters correctly; however, some are not.
4: The response uses most of the parameters correctly except a small number of minor errors.
5: The response uses all parameters correctly.
(7) Readability: the easiness for human readers to read and understand the produced solution. This is assessed by a human tester to read the produced solution and attempt to understand it, where a solution typically includes pieces of code together with natural language explanations. The scores are:
1: The response is difficult to understand, with poor grammar, confusing language and awkward spacing, etc.
2: The response is somewhat readable but contains poor grammar, confusing language, and some awkward format, such as spacing.
3: The response is mostly readable, but it includes some grammar issues, confusing language and awkward spacing.
4: The response is generally well-written and easy to understand, with only minor grammar, language, and formatting issues.
5: The response is highly readable, language is clear and concise, grammar is correct, formatting is good.
(8) Depth of explanation: the quality of the textual explanation produced together with the code. This is assessed by reviewing the text produced by the language model accompanying the code and provides explanation of the generated code. It focuses on how well the text explains the code in a comprehensible and understandable way and to elucidate the rationale behind the code. The scores are:
1: The text part of the response is in lack of information and explanation of the code.
2: The text part of the response provides some basic information about the code, but many aspects are not fully explained.
3: The text part of the response provides the necessary information and explanation of the code; however, the explanation lacks depth.
4: The text part of the response provides sufficient explanations on most aspects of the code at a satisfactory level of detail and depth.
5: The text part of the response produces highly sufficient explanations that covers all aspects of the code with high clarity and depth.

### v3

@@ V-C Impacts of Difficulty Levels on Performance
Moreover, when measuring usability on each individual quality criterion, the experiment data also clearly demonstrate the monotonicity of performance change with the difficulty level; see Figure 13.
@@
Figure 13: Usability on Various Quality Criteria on Different Difficulty Levels
V-D Potential Threats to The Validity of Experiment Results
@@
The scale of our experiments is small in terms of the number of subjects involved in the testing process due to our resource limitation. This forms a potential threat to the validity and generalizability of the results. However, the feasibility of the proposed testing method is not affected. The potential threat can be eliminated by repeated experiments with more subjects. This suggests a topic for future work to repeat the experiments with more testers.
@@
The manual assessment on quality attributes is subjective, which depends on the subject’s knowledge, experience, and ability. This causes another potential threat to the validity of the results. Efforts have already been made to confine the threat by applying the metrics consistently according to the clear definitions of the metric values. Thus, the risk is minimized. Moreover, our experiment data show that the results obtained from such subjective judgements are consistent with objective metrics of the number of attempts and time to completion. Thus, the results should be trustworthy. For future research, we are working on using more objective metrics about the quality of program code developed in software engineering research, like cyclomatic complexity, coupling and cohesion, code smell, etc., in complementary to the subjective metrics proposed and used in this paper. It is worth further research to find out if machine generated codes are of high quality on these attributes designed for traditional human written code, and whether such metrics can provide meaningful assessment of the usability of machine generated code.
@@
A factor that has an impact on the experiment results is the criterion to determine whether a solution is satisfactory in the multi-attempt testing process. In principle, being user centric, it must be judged by the human users who engage in the testing. In our experiment, the guideline given to the tester is that the generated code is satisfactory if it is correct when compared with the reference solution or is close enough to the reference solution. In other words, a functionally incorrect solution may well be regarded as satisfactory if a minor editing of the machine generated solution will work. However, in practice, such reference solutions are usually unknown. Therefore, there is a gap between the judgements of satisfactoriness by the testers and the real users. This is a potential threat to the validity of the experiment results. It is desirable to give a more rigorous criterion to judge how satisfactory a generated solution is. However, defining such a rigorous criterion is a non-trivial issue and worth further research. Study of metrics on the usefulness of machine generated solutions could be helpful.
@@
Our dataset is constructed by extracting exercise questions from textbooks. An advantage is that the questions are well presented in terms of the clarity of the language and coverage of the programming language facilities and usages as well as the correctness of the reference solution. Thus, little manual editing and filtering of data is required while other benchmarks constructed by extracting data from open-source code repositories have required heavy workload on manual editing and cleaning of the data. Consequently, we have managed to construct a relatively large dataset with a small amount of human resource. It also minimizes the possibility of testing LLM on incorrect questions. A potential threat to the validity of the experiment is that questions from textbooks may not represent the coding tasks in the real uses of LLMs. It is worth further research on the validity of the dataset via studying the differences between real input from users and the data in our benchmark.
@@
We have included three types of metadata to our benchmark data: the type of code to be generated, the difficulty level of the task to be completed, and the source of the data. The first two represent two kinds of scenarios, which can be combined to form various specific usage scenarios as demonstrated in the case study. Since the metadata are manually assigned to the test cases in the dataset, a potential threat to the validity of the experiment results can be caused by errors in the metadata. In the experiment, we have taken mitigation means to minimize the error rate by giving clear definitions of the values of metadata. For example, a coding task is assigned with “hard” as its difficulty level, if and only if “when the program code to be generated requires complex data structures and algorithms.” In the literature, some datasets do classify the coding tasks into different difficulty levels. For example, APPS dataset has three subsets of data on three difficult levels: Introductory, Interview, and Competition. However, there is no definition of the criterion on how the tasks were classified [24]. Our approach enables metadata to be assigned systematically and consistently, thus the chances of human errors are reduced. As for future work, it is worth studying other types of metadata and other values of metadata, for example, difficulty levels could be more than three. Another topic for future work is how to check the correctness of metadata.
VI Comparison with Existing Works
Existing evaluations of LLM models are mostly conducted through benchmarking on given test datasets. While this enables objective comparisons of LLMs, the results hardly reflect their usability from the user’s perspective. The method proposed in this paper is a scenario-based approach. It aims at evaluation of a LLM on a number of scenarios that each scenario is represented by a subset of test cases. Understanding the performances of a LLM on various different scenarios is particularly important when deciding on whether to use a LLM in a specific use scenario. The difficulty is that there is a wide range potential uses of LLMs, while testing and evaluating a LLM is expensive and difficult. Our solution is to assign metadata to each test case in the benchmark dataset to describe various aspects of its usages. This enables flexible tailoring of the dataset to generate subsets that represent various use scenarios, and to adjust the evaluation on different scenarios with existing test results without re-run test cases. It is different from existing benchmarks for evaluating code generation capability. As discussed in Section II, these benchmarks only contain the coding task, the reference solution(s) and/or a set of test data for checking the correctness of the machine generated solution.
Our method also differs from existing work on the testing process. Our multi-attempt process of testing on each test case mimics the real uses of LLMs. In this process, the tester tries to complete the testing task through a series of attempts until a satisfactory solution is obtained or gives up after a fixed number of maximal allowed attempts. In each attempt, the tester formulates an input based on the previous results and the task to be completed, invokes the ML model, and then inspects the output to decide whether the solution is satisfactory. As discussed in Section II, this contrasts with the multi-trial/single-attempt process currently used by the existing work on evaluating code generation. In the multi-trial process, a number of solutions are generated for the same input. Our process is closer to the real uses of LLMs.
Moreover, in our method, the quality of an output is measured by a set of metrics on the usefulness of the generated code from the users’ perspective. This contrasts with existing works on evaluations of code generation, where the functional equivalence of a machine produced solution to the reference solution(s) are used as the sole judgement on the quality of the code generated. In the real uses of LLMs for code generation, a piece of code is considered as useful even if it is not functionally correct. In many cases, a skeleton of code is acceptable even if it cannot be compiled. Our quality attributes focus on the usability of a solution via considering its readability, logic clarity, well-structuredness, etc. Thus, again the method is closer to the real uses of LLMs.
Finally, we define two metrics to measure user experiences in the uses of a LLM: (1) #$\#attempt_{k}$, i.e. the average number of attempts, and (2) the average task complete time. They differ from the $pass@k$ metric, which is currently widely used in evaluation of code generation capability. Our metrics better reflect the user experiences in the use of the LLM.
Therefore, our proposed approach to the evaluation of LLM on code generation can be characterized as user centric for its focus on user experiences in the uses of the LLM. This is reflected in all aspects of the testing and evaluation, including the information contained in the benchmark dataset, the testing process, the measurement of solution quality and the performance metrics.
Our case study results are consistent with the good user experiences and general opinions on ChatGPT. We have also identified the weakness as well as the strength of the LLM model.
VII Conclusion
This paper proposed a user centric methodology to the testing and evaluation of ML models for code generation. The main contributions are as follows.
First, we introduced the uses of metadata assigned to test cases in a benchmark to support flexible construction of test datasets that represent various usage scenarios of ML models and flexible analysis of testing results.
Second, we proposed the multi-attempt testing process so that the testing on each test case mimics the real uses of a ML model by the user.
Third, we defined a new the performance metric #$\#Attempt_{k}$, i.e. the average number of attempts . Together with the average task completion time, it provides an objective measure of the user experiences in the use of a ML model to generate program code.
Finally, we employed a set of quality attributes to measure the usability of generated program code and explanation text.
The proposed methodology covers all aspects of ML model evaluation from the construction of benchmark dataset, testing process, quality attributes to be assessed, to the metrics to evaluate ML model’s performance. In comparison with the existing approaches, its main advantage is that the evaluation results can better reflect users’ experiences in the use of a ML model. Moreover, by providing evaluation of a ML model on a spectrum of scenarios, the results can provide more insights of the weakness and strengths of the model for developers to improve its performance.
@@ VII Conclusion
There are several interesting topics worth research to further develop the methodology. For example, the user centric testing method proposed in this paper is expensive to apply because of human users’ involvement in the testing process. It is worth further research to reduce the cost and to improve efficiency through testing tools and automation.
@@
It is worth noting that evaluating LLM is inevitably subjective due to the nature of NLP. Instead of trying to completely avoid subjective assessment, we minimize the potential bias of the assessment by decomposing the overall subject opinions into a set of quality attributes that represent different aspects of subjective opinions and use Likert’s 5 levels of scores to normalize the assessments. This approach has been well developed and widely applied in software engineering and HCI to evaluate software quality. How to further improve the methodological rigour of user centric evaluations of LLMs will be of great importance.
@@
Finally, a LLM’s capability of generating program code is not only useful to software development as a coding tool, but also for many other purposes such as education. For example, a student could learn coding with a LLM by raising questions in natural language and getting answers from the LLM with program codes and detailed explanations. Another possibility is to help students to understand example program code by inputting a piece of code to a LLM and getting an explanation of the code. Moreover, a student’s answer to an exercise or examination question could be checked and marked by a LLM for its correctness and producing a feedback to the student. If a student’s answer to an exercise question is incorrect, the LLM could give the reasons why it is incorrect and delineate on how to correct it. It is worth further research to develop a methodology for the evaluation of LLMs in such uses based on the user centric principle.